Volume of oxygen at ntp required to completely burn 2 kg of coal (100%carbon)is

Grade:11

1 Answers

Arun
5 years ago
equation: 
C + O2 → CO2 
1mol C will react with 1 mol O2 
Mass ratio 12:32 
2 kg coal will require 2 *32/12 = 5.334kg = 5334 g O2 
Molar mass O2 = 32g/mol 
5334g = 5334/32 = 166.68mol 
Ar NTP 1 mol = 24L 
166.68 mol = 166.68*24 = 4000 L 
You ask about NTP = my understanding is that at NTP 1 mol gas = 24L 

But at STP , 1 mol gas = 22.4L 
At STP , 166.68 mol = 166.68 *22.4 = 3534 L OR 3.534*10^3 L 

Volume of O2 required art STP = 3.534*10^3 L